# Convergence order of the quantization error for self-affine measures on Lalley-Gatzouras carpets

### v1

@@ 3.2. Auxiliary measures
Using Lemmas 3.4, 3.8, we have
$I_{k+p,r}(t)=\sum_{\sigma\in\Phi_{k}}\sum_{\rho\in\Lambda_{p}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\geq f_{1}(t)I_{k,r}(t)I_{p,r}(t).$
@@
As a consequence of Lemma 3.9, we can obtain the following standard result.
@@
(1) For every $t>0$, $\lim_{k\to\infty}\frac{1}{k}\log I_{k,r}(t)=:g(t)$ exists. (2) There exists a unique $s_{r}>0$ such that for $t_{r}=\frac{s_{r}}{s_{r}+r}$, we have $g(t_{r})=0$. (3) For $C(t):=f_{1}(t)^{-1}f_{2}(t)$ and $k,p\geq 1$, we have,
$C(t)^{-1}I_{k}(t_{r})\leq I_{p}(t_{r})\leq C(t)I_{k}(t_{r}).$
@@
This can be proved by using Lemma 3.9, the fact that $\dim_{B}E>0$ and [5, Corollary 1.2] and along the line of [5, Lemma 5.2]. ∎
With the help of Lemmas 3.4-3.9, we are now able to construct an auxiliary probability measure on $\Phi_{\infty}$ by applying Prohorov’s theorem. Recall that a family $\mathcal{\pi}$ of probability measures on a metric space $X$, is said to be tight if for every $\epsilon\in(0,1)$, there exists some compact subset $K$ of $X$ such that $\inf_{\nu\in\mathcal{\pi}}\nu(K)\geq 1-\epsilon$. In particular, if $X$ is a compact metric space, then every family $\mathcal{\pi}$ of probability measures on $X$ is tight.
@@
There exists a Borel probability measure $\lambda$ on $\Phi_{\infty}$ such that, for every $\sigma\in\Phi^{*}$, we have $\lambda([\sigma])\asymp\mathcal{E}_{r}(\sigma)^{t_{r}}$.
@@
Then $(\lambda_{k})_{k=1}^{\infty}$ is a sequence of probability measures on $\Phi_{\infty}$. Note that $\Phi_{\infty}$ is a compact metric space; so $(\lambda_{k})_{k=1}^{\infty}$ is tight. According to Prohorov’s Theorem (cf. [3, Theorem 5.1]), there exists a subsequence $(\lambda_{k_{i}})_{i=1}^{\infty}$ and a Probability measure $\lambda$ on $\Phi_{\infty}$ such that $\lambda_{k_{i}}$ converges weakly to $\lambda$. Let $k\geq 1$ and $\sigma\in\Phi_{k}$ be given. For every $p>k$, using Lemmas 3.4-3.9, we deduce
$\lambda_{p}([\sigma])$ $=$ $\sum_{\rho\in\Lambda_{p-k}(\sigma)}\lambda_{p}([\rho])=\frac{1}{I_{p,r}(t_{r})}\sum_{\rho\in\Lambda_{p-k}(\sigma)}\mathcal{E}_{r}(\rho)^{t_{r}}$
$\asymp$ $\frac{1}{I_{p,r}(t_{r})}\mathcal{E}_{r}(\sigma)^{t_{r}}I_{p-k,r}(t_{r})\asymp\mathcal{E}_{r}(\sigma)^{t_{r}}.$
Note that $\sigma_{L}]$ ($\sigma_{R}]$) is both open and closed in $G^{\mathbb{N}}$ ($G_{y}^{\mathbb{N}}$). Thus $\sigma]$ is clopen in $\Phi_{\infty}$ (for closeness, it suffices to note that its complement is open). Because $\Phi_{\infty}$ is compact, we deduce that $\sigma]$ is also compact. It follows that $\lambda([\sigma])\asymp\mathcal{E}_{r}(\sigma)^{t_{r}}$. ∎
4. Proof of Theorem 1.2
By [31, Lemma 3.4], for the proof of Theorem 1.2, it is sufficient to show
@@
Next, we give the proof for the first inequality of (4.1), in an analogous manner to that for [30, Proposition 3.2].
@@
By Lemma 3.1, for $\hat{\sigma},\hat{\omega}\in\hat{\Lambda}_{n,r}$, the sets $\hat{\sigma}],[\hat{\omega}$], are either disjoint, or one is contained in the other. Also, by Remark 2.7 (1), we have
@@
If $\hat{\omega}]\subset[\hat{\sigma}$], then as we did in [30, Lemma 3.1], one can see that, there exists some constant $H_{1,r}\geq 1$ such that $\big{|}|\hat{\sigma}_{L}|-|\hat{\omega}_{L}|\big{|}\leq H_{1,r}$. This allows us to select a subset $\hat{\Lambda}_{n,r}^{\flat}$ of $\hat{\Lambda}_{n,r}$ such that, the sets $\hat{\sigma}],\hat{\sigma}\in\hat{\Lambda}_{n,r}^{\flat}$, are pairwise disjoint, and
$\hat{\Lambda}_{n,r}=\bigcup_{\hat{\sigma}\in\hat{\Lambda}_{n,r}^{\flat}}\Gamma(\hat{\sigma}),\;{\rm with}\;\;\Gamma(\hat{\sigma})=\{\mathcal{L}(\omega):\hat{\sigma}\preceq\mathcal{L}(\omega),\omega\in\Lambda_{n,r}\}.$
@@
$\sum_{\hat{\sigma}\in\hat{\Lambda}_{n,r}}\mathcal{E}_{r}(\hat{\sigma})^{t_{r}}\asymp\sum_{\sigma\in\hat{\Lambda}_{n,r}}\lambda([\hat{\sigma}])\leq H_{1,r}\sum_{\sigma\in\hat{\Lambda}_{n,r}^{\flat}}\lambda([\hat{\sigma}])\leq H_{1,r}.$
@@
In the following, we are going to prove the last inequality in (4.1). We need to define the predecessors for $\sigma\in\Phi^{*}$. For $\sigma\in\Phi_{1}$, we define $\sigma^{-}:=\theta$. Let $l\geq 2$ and $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\times(j_{l+1},\ldots,j_{k})\in\Phi_{l}$, we have, $b_{\sigma_{y}^{-}}\geq a_{\sigma_{L}}>b_{\sigma_{y}}$. Hence,
@@
There exists a unique integer $h\geq 0$, such that the following inequalities hold:
$\prod_{h=1}^{l-1}b_{j_{h}}\prod_{h=l+1}^{k-h-1}b_{j_{h}}\geq\prod_{h=1}^{l-1}a_{i_{h}j_{h}}>\prod_{h=1}^{l-1}b_{j_{h}}\prod_{h=l+1}^{k-h}b_{j_{h}}.$
For this integer $h$, we define $\sigma^{-}:=\sigma_{L}^{-}\times(j_{l+1},\ldots,j_{k-h})$. By Lemma 3.6, we obtain
(4.2) $\underline{\eta}_{r}\mathcal{E}_{r}(\sigma^{-})\leq\mathcal{E}_{r}(\sigma)\leq\overline{p}\;\overline{a}^{r}\mathcal{E}_{r}(\sigma)<\overline{\eta}_{r}\mathcal{E}_{r}(\sigma^{-}).$
@@
This contradicts the definition of $G_{1}(\sigma)$. If follows that $G_{1}(\sigma)\subset\bigcup_{h=1}^{M_{r}}\Gamma_{h}(\sigma)$, where
(4.4) $\Gamma_{h}(\sigma):=\{\rho\in\Psi_{l+h}:\sigma\preceq\rho\}.$
The following lemma is an analogue of [32, Lemma 4.1].
@@
By Lemma 2.4 (i) and a coarse estimate, for every $h\geq 1$, we have that ${\rm card}(\Gamma_{h}(\sigma))\leq N^{h}m^{hA_{1}}$. This, along with (2.1), yields
$\sum_{\omega\in G_{1}(\sigma)}\mathcal{E}_{r}(\omega)^{t_{r}}\leq\sum_{h=1}^{M_{r}}\sum_{\omega\in\Gamma_{h}(\sigma)}\mathcal{E}_{r}(\omega)^{t_{r}}\leq\sum_{h=1}^{M_{r}}{\rm card}(\Gamma_{h}(\sigma))\overline{\eta}_{r}^{ht_{r}}\mathcal{E}_{r}(\sigma)^{t_{r}}.$
It is sufficient to define $H_{2,r}:=\sum_{h=1}^{M_{r}}N^{h}m^{hA_{1}}\overline{\eta}_{r}^{ht_{r}}$. ∎
@@
From (4.2) and the definition of $\Gamma_{n,r}$, we know that, $\widetilde{\Gamma}_{n,r}\subset\mathcal{S}_{n,r}$; and every $\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}$ corresponds to an approximate square. In view of Remark 3.2 (r2), we define
@@
$\Lambda_{n,r}^{\flat}:=\{\sigma\in\Lambda_{n,r}:F^{\circ}_{\sigma}\cap\bigg{(}\bigcup_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}}F_{\widetilde{\sigma}}^{\circ}\bigg{)}\neq\emptyset\}$.
@@
$\widetilde{\Gamma}_{n,r}^{\flat}(1):=\{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}:F_{\sigma}\subseteq F_{\widetilde{\sigma}}\;{\rm for\;some}\;\sigma\in\Lambda_{n,r}\},$
$\widetilde{\Gamma}_{n,r}^{\flat}(2):=\{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}:F_{\sigma}\supsetneq F_{\widetilde{\sigma}}\;{\rm for\;some}\;\sigma\in\Lambda_{n,r}\}.$
By Lemma 3.1, $\Lambda_{n,r}^{\flat}$ is the disjoint union of $\Lambda_{n,r}^{\flat}(1)$ and $\Lambda_{n,r}^{\flat}(2)$. We further define
@@
$A_{1,r}:=\sum_{\sigma\in\Lambda_{n,r}^{\flat}(1)}\mathcal{E}_{r}(\sigma)^{t_{r}};\;A_{2,r}:=\sum_{\sigma\in\Lambda_{n,r}^{\flat}(2)}\mathcal{E}_{r}(\sigma)^{t_{r}}.$
In the following, we estimate $A_{1,r}$ and $A_{2,r}$ separately.
@@
$\sum_{\sigma\in T(\widetilde{\sigma})}\mathcal{E}_{r}(\sigma)^{t_{r}}=\sum_{\sigma\in T(\widetilde{\sigma})}(\mu(F_{\sigma})a_{\sigma_{L}^{r}})^{t_{r}}\geq\underline{a}^{\frac{M_{r}rs_{r}}{s_{r}+r}}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}.$
Let $H_{3,r}:=\underline{a}^{\frac{M_{r}rs_{r}}{s_{r}+r}}$. It follows that
(4.7) $A_{1,r}=\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}(1)}\sum_{\sigma\in T(\widetilde{\sigma})}\mathcal{E}_{r}(\sigma)^{t_{r}}\geq H_{3,r}\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}(1)}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}.$
For every $\sigma\in\Lambda_{n,r}^{\flat}(2)$, we have, $S(\sigma)\subseteq G_{1}(\sigma)$. Thus, by Lemma 4.3, we deduce
(4.8) $A_{2,r}\geq\sum_{\sigma\in\Lambda_{n,r}^{\flat}(2)}H_{2,r}^{-1}\sum_{\widetilde{\sigma}\in S(\sigma)}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}=H_{2,r}^{-1}\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}(2)}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}$

### v10

@@ 3.2. Auxiliary measures
For $\sigma\in\Phi^{*}$ and $h\geq 1$, we have ${\rm card}(\Lambda_{h}(\sigma))\geq N^{h}$. By Lemma 3.4,
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\geq N^{h}\underline{\eta}_{r}^{ht}\mathcal{E}_{r}(\sigma)^{t}.$
Let $A_{7}:=[\frac{\log\underline{b}}{\log\overline{b}}]+2$ and $A_{8}:=[A_{7}A_{4}^{-1}]+1$. We define
$\xi_{2,r}(t):=\min\limits_{1\leq h\leq A_{8}}(I_{h,r}(t)^{-1}N^{h}\underline{\eta}_{r}^{ht}).$
Then for every $1\leq h\leq A_{8}$, we have
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\geq\xi_{2,r}(t)\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
[…]
It follows that
$\prod_{p=1}^{k-1}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}}b_{\hat{j}_{p}}.$
We need to distinguish between the following two cases:
  1. (1)
    $\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}$. In this case, we define
    We have that $\rho(\omega)\in\Lambda_{h}(\sigma)$. We denote by $\widehat{\Lambda}_{h,1}(\sigma)$ the of set such words $\rho(\omega)$ and denote the set of the words $\omega$ by $\widehat{\Phi}_{h,1}$.
  2. (2)
    $\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}<\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}$. We have
    (3.6) $\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-A_{7}}b_{\hat{j}_{p}}$ $=$ $\prod_{p=1}^{k-1}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\frac{b_{j_{k}}}{\prod_{p=\hat{k}-A_{7}+1}^{\hat{k}-1}b_{\hat{j}_{p}}}$
    $\geq$ $\prod_{p=1}^{k-1}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\frac{\underline{b}}{\overline{b}^{A_{7}-1}}$
    $>$ $\prod_{p=1}^{k-1}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}.$
    Hence, there exists a unique integer $2\leq H\leq A_{7}$, such that
    $\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-H}b_{\hat{j}_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-H+1}b_{\hat{j}_{p}}.$
    Let $\omega_{R}|_{\hat{k}-h-H+1}:=(\hat{j}_{h+1},\ldots,\hat{j}_{\hat{k}-H+1})$. We define
    $\rho(\omega):=(\sigma_{L}\ast\omega_{L})\times(\sigma_{R}\ast(\omega_{R}|_{\hat{k}-h-H+1}).$
    Then $\rho(\omega)\in\Lambda_{h}(\sigma)$. We denote the set of such words $\rho(\omega)$ by $\widehat{\Lambda}_{h,2}(\sigma)$ and denote the set of the words $\omega$ by $\widehat{\Phi}_{h,2}$.
We need to observe the following facts.
(i) We clearly have that, $\widehat{\Phi}_{h,1}\cap\widehat{\Phi}_{h,2}=\emptyset,\Phi_{h}=\widehat{\Phi}_{h,1}\cup\widehat{\Phi}_{h,2}$. Also, we have, $\widehat{\Lambda}_{h,1}(\sigma)\cap\widehat{\Lambda}_{h,2}(\sigma)=\emptyset$. In fact, a word $\rho(\omega)\in\widehat{\Lambda}_{h,1}(\sigma)$ can not be obtained by any $\tau\in\widehat{\Phi}_{h,2}$ and vice versa. Otherwise, we would have $\omega_{L}=\tau_{L}$ and $\tau_{R}\precneqq\omega_{R}$, contradicting the fact that $\omega,\tau\in\Phi_{h}$.
(ii) For different words $\omega,\tau\in\widehat{\Phi}_{h,1}$, we have $\rho(\omega)\neq\rho(\tau)$.
(iii) There exist at most $\sum_{p=1}^{A_{7}}m^{h}$ words in $\widehat{\Phi}_{h,2}$ that determine the same word in $\widehat{\Lambda}_{h,2}(\sigma)$, because of the absence of $\hat{j}_{\hat{p}},\hat{k}-H+2\leq p\leq\hat{k}$ in $\rho(\omega)$. To see this, we fix an $\omega\in\widehat{\Phi}_{h,2}$. For $\tau\in G_{y}^{*}$, let $\hat{\omega}^{(\tau)}:=\omega_{L}\times(\omega_{R}|_{\hat{k}-h-H+1}\ast\tau)$. Let $\rho^{-1}(\rho(\omega)):=\{\hat{\omega}\in\Phi_{h,2}:\rho(\hat{\omega})=\rho(\omega)\}$. We have
$\rho^{-1}(\rho(\omega)):=\langle\omega\rangle=\big\{\hat{\omega}^{(\tau)}:\tau\in G_{y}^{*}\big\}\cap\Phi_{h,2}.$
For every $\hat{\omega}^{(\tau)}\in\langle\omega\rangle$, by (3.6), we have $|\tau|<A_{7}$. It follows that
(3.7) $\langle\omega\rangle\subset\bigg\{\hat{\omega}^{(\tau)}:\tau\in\bigcup_{h=1}^{A_{7}}G_{y}^{h}\bigg\}.$
For $\omega\in\widehat{\Phi}_{h,2}$, we take an arbitrary word of $\langle\omega\rangle$ and denote the set of such words by $\widehat{\Phi}^{\flat}_{h,2}$. By (3.7),
(3.8) $\sum_{\omega\in\widehat{\Phi}_{h,2}}\mathcal{E}_{r}(\omega)^{t}=\sum_{\omega\in\widehat{\Phi}_{h,2}^{\flat}}\sum_{\widetilde{\omega}\in\langle\omega\rangle}\mathcal{E}_{r}(\widetilde{\omega})^{t}\leq\sum_{\omega\in\widehat{\Phi}_{h,2}^{\flat}}\frac{\mathcal{E}_{r}(\rho(\omega))^{t}}{\mathcal{E}_{r}(\sigma)^{t}}\sum_{p=1}^{A_{7}}\big(\sum_{j\in G_{y}}q_{j}^{t}\big)^{p}.$
Let $\xi_{3}(t):=\sum_{p=1}^{A_{7}}\big(\sum_{j\in G_{y}}q_{j}^{t}\big)^{p}$. Using (3.8), we deduce
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}$ $\geq$ $\sum_{\rho\in\widehat{\Lambda}_{h,1}(\sigma)}\mathcal{E}_{r}(\rho)^{t}+\sum_{\rho\in\widehat{\Lambda}_{h,2}(\sigma)}\mathcal{E}_{r}(\rho)^{t}$
$=$ $\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\widehat{\Phi}_{h,1}}\mathcal{E}_{r}(\omega)^{t}+\sum_{\omega\in\widehat{\Phi}^{\flat}_{h,2}}\mathcal{E}_{r}(\rho(\omega))^{t}$
$\geq$ $\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\widehat{\Phi}_{h,1}}\mathcal{E}_{r}(\omega)^{t}+\mathcal{E}_{r}(\sigma)^{t}\xi_{3}(t)^{-1}\sum_{\omega\in\widehat{\Phi}_{h,2}}\mathcal{E}_{r}(\omega)^{t}$
$\geq$ $\min\bigg\{1,\xi_{3}(t)^{-1}\bigg\}\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
Thus, (3.4) is fulfilled by defining $h_{2,r}(t):=\min\{1,\xi_{3}(t)^{-1},\xi_{2,r}(t)\}$. ∎
Combining Lemmas 3.5, 3.6, we obtain
Lemma 3.7.
For every $t\geq 0$ and $k,p\in\mathbb{N}$, we have
$h_{2,r}(t)I_{k,r}(t)I_{p,r}(t)\leq I_{k+p,r}(t)\leq h_{1,r}(t)I_{k,r}(t)I_{p,r}(t).$
Proof.
Using Lemmas 3.5, 3.6, we have
$I_{k+p,r}(t)=\sum_{\sigma\in\Phi_{k}}\sum_{\rho\in\Lambda_{p}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\geq h_{2,r}(t)I_{k,r}(t)I_{p,r}(t).$
@@
Using Lemma 3.7, we can obtain the following standard result.
@@
(1) For every $t\geq 0$, $\lim_{k\to\infty}\frac{1}{k}\log I_{k,r}(t)=:g_{r}(t)$ exists. (2) There exists a unique $s_{r}>0$ such that for $t_{r}=\frac{s_{r}}{s_{r}+r}$, we have $g_{r}(t_{r})=0$. (3) For $C(t):=h_{2,r}(t)^{-1}h_{1,r}(t)$ and $k,p\geq 1$, we have,
$C(t)^{-1}I_{k,r}(t_{r})\leq I_{p,r}(t_{r})\leq C(t)I_{k,r}(t_{r}).$
@@
(1) is an easy consequence of Lemma 3.7 and [3, Corollary 1.2]. (2) Along the line of [3, Lemma 5.2], one can see that $g_{r}$ is strictly decreasing and continuous. Note that $g_{r}(0)\geq\log N>0,g_{r}(1)\leq r\log\overline{a}<0$. Thus, (2) follows from the continuity of $g_{r}$. (3) For every $k\geq 1$, we have (cf. [3, (5.10)])
$\frac{1}{k}(\log I_{k,r}(t_{r})+\log h_{1,r}(t_{r}))\geq\inf_{p\geq 1}\frac{1}{p}(\log I_{p,r}(t_{r})+\log h_{1,r}(t_{r}))=0.$
It follows that $I_{k,r}(t_{r})\geq h_{1,r}(t_{r})^{-1}$. Similarly, we have $I_{k,r}(t_{r})\leq h_{2,r}(t_{r})^{-1}$. This completes the proof of (3). ∎
With the help of Lemmas 3.5, 3.6 and Proposition 3.8, we are now able to construct an auxiliary probability measure on $\Phi_{\infty}$ by applying Prohorov’s theorem. Recall that a family $\mathcal{\pi}$ of probability measures on a metric space $X$, is said to be tight if for every $\epsilon\in(0,1)$, there exists some compact subset $K$ of $X$ such that $\inf_{\nu\in\mathcal{\pi}}\nu(K)\geq 1-\epsilon$. In particular, if $X$ is a compact metric space, then every family $\mathcal{\pi}$ of probability measures on $X$ is tight.
@@
There exists a Borel probability measure $\lambda$ on $\Phi_{\infty}$ such that
$\lambda([\sigma])\asymp\mathcal{E}_{r}(\sigma)^{t_{r}},\;\sigma\in\Phi^{*}.$
@@
Note that $\Phi_{\infty}$ is a compact metric space; so $(\lambda_{k})_{k=1}^{\infty}$ is tight. By Prohorov’s Theorem (cf. [2, Theorem 5.1]), there exist a subsequence $(\lambda_{k_{i}})_{i=1}^{\infty}$ and a probability measure $\lambda$ on $\Phi_{\infty}$ such that $\lambda_{k_{i}}$ converges weakly to $\lambda$. Let $n\geq 1$ and $\sigma\in\Phi_{n}$ be given. For every $i>n$, using Lemmas 3.5, 3.6 and Proposition 3.8, we deduce
(3.9) $\lambda_{k_{i}}([\sigma])=\sum_{\rho\in\Lambda_{k_{i}-n}(\sigma)}\lambda_{k_{i}}([\rho])=\frac{1}{I_{k_{i},r}(t_{r})}\sum_{\rho\in\Lambda_{k_{i}-n}(\sigma)}\mathcal{E}_{r}(\rho)^{t_{r}}\asymp\mathcal{E}_{r}(\sigma)^{t_{r}}.$
Note that $\sigma_{L}]$ ($\sigma_{R}]$) is both open and closed in $G^{\mathbb{N}}$ ($G_{y}^{\mathbb{N}}$). Thus $\sigma]$ is clopen in $\Phi_{\infty}$. Because $\Phi_{\infty}$ is compact, $\sigma]$ is also compact. By (3.9), it follows that $\lambda([\sigma])\asymp\mathcal{E}_{r}(\sigma)^{t_{r}}$. ∎
4. Proofs of Theorem 1.2 and Proposition 1.4
4.1. Proof of Theorems 1.2
By [27, Lemma 3.4], for the proof of Theorem 1.2, it is sufficient to show
@@
Next, we give the proof for the first inequality of (4.1), in an analogous manner to that for [26, Proposition 3.2].
@@
By Lemma 3.1, for $\hat{\sigma},\hat{\omega}\in\hat{\Lambda}_{n,r}$, either the sets $\hat{\sigma}],[\hat{\omega}$], are disjoint, or one is contained in the other. Also, by Remark 2.4 (1), we have
@@
If $\hat{\omega}]\subset[\hat{\sigma}$], then as we did in [26, Lemma 3.1], one can see that, there exists a constant $H_{1,r}\geq 1$ such that $\big||\hat{\sigma}_{L}|-|\hat{\omega}_{L}|\big|\leq H_{1,r}$. This allows us to select a subset $\hat{\Lambda}_{n,r}^{\flat}$ of $\hat{\Lambda}_{n,r}$ such that, the sets $\hat{\sigma}],\hat{\sigma}\in\hat{\Lambda}_{n,r}^{\flat}$, are pairwise disjoint and
$\hat{\Lambda}_{n,r}=\bigcup_{\hat{\sigma}\in\hat{\Lambda}_{n,r}^{\flat}}\Gamma(\hat{\sigma}),\;{\rm with}\;\;\Gamma(\hat{\sigma})=\{\hat{\omega}:\hat{\sigma}\preceq\hat{\omega},\hat{\omega}\in\hat{\Lambda}_{n,r}\}.$
@@
$\sum_{\hat{\sigma}\in\hat{\Lambda}_{n,r}}\mathcal{E}_{r}(\hat{\sigma})^{t_{r}}\asymp\sum_{\sigma\in\hat{\Lambda}_{n,r}}\lambda([\hat{\sigma}])\leq(H_{1,r}+1)\sum_{\sigma\in\hat{\Lambda}_{n,r}^{\flat}}\lambda([\hat{\sigma}])\leq H_{1,r}+1.$
@@
In the following, we are going to prove the last inequality in (4.1). For every $\sigma\in\Phi_{1}$, we define $\sigma^{-}:=\theta$. Let $l\geq 2$ and $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\times(j_{l+1},\ldots,j_{k})\in\Phi_{l}$, we have,
@@
There exists a unique integer $p\geq 0$, such that the following inequalities hold:
$\prod_{h=1}^{l-1}b_{j_{h}}\prod_{h=l+1}^{k-p-1}b_{j_{h}}\geq\prod_{h=1}^{l-1}a_{i_{h}j_{h}}>\prod_{h=1}^{l-1}b_{j_{h}}\prod_{h=l+1}^{k-p}b_{j_{h}}.$
We define $\sigma^{-}:=\sigma_{L}^{-}\times(j_{l+1},\ldots,j_{k-p})$. By Lemma 3.4, we obtain
(4.2) $\underline{\eta}_{r}\mathcal{E}_{r}(\sigma^{-})\leq\mathcal{E}_{r}(\sigma)\leq\overline{p}\;\overline{a}^{r}\mathcal{E}_{r}(\sigma^{-})<\overline{\eta}_{r}\mathcal{E}_{r}(\sigma^{-}).$
@@
This contradicts the definition of $G_{1}(\sigma)$. Thus, for $\sigma\in\mathcal{S}_{n,r}\cap\Psi_{l}$, we have that $G_{1}(\sigma)\subset\bigcup_{h=0}^{M_{r}}\Gamma_{h}(\sigma)$, where $\Gamma_{h}(\sigma):=\{\rho\in\Psi_{l+h}:\sigma\preceq\rho\}$.
The following lemma is an analogue of [28, Lemma 4.1].
@@
By Lemma 2.2 (i), for every $h\geq 1$, we have that ${\rm card}(\Gamma_{h}(\sigma))\leq N^{h}m^{hA_{1}}$. This, along with (2.4), yields
$\sum_{\omega\in G_{1}(\sigma)}\mathcal{E}_{r}(\omega)^{t_{r}}\leq\sum_{h=0}^{M_{r}}\sum_{\omega\in\Gamma_{h}(\sigma)}\mathcal{E}_{r}(\omega)^{t_{r}}\leq\sum_{h=0}^{M_{r}}N^{h}m^{hA_{1}}\overline{\eta}_{r}^{ht_{r}}\mathcal{E}_{r}(\sigma)^{t_{r}}.$
It is sufficient to define $H_{2,r}:=\sum_{h=0}^{M_{r}}N^{h}m^{hA_{1}}\overline{\eta}_{r}^{ht_{r}}$. ∎
@@
From (4.2) and the definition of $\Gamma_{n,r}$, we know that, $\widetilde{\Gamma}_{n,r}\subset\mathcal{S}_{n,r}$; and every $\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}$ corresponds to an approximate square. We define
@@
[…]
@@
$\widetilde{\Gamma}_{n,r}^{\flat}(1):=\{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}:F_{\sigma}\subseteq F_{\widetilde{\sigma}}\;{\rm for\;some}\;\sigma\in\Lambda_{n,r}^{\flat}\},$
$\widetilde{\Gamma}_{n,r}^{\flat}(2):=\{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}:F_{\sigma}\supsetneq F_{\widetilde{\sigma}}\;{\rm for\;some}\;\sigma\in\Lambda_{n,r}^{\flat}\}.$
By Lemma 3.1, $\Lambda_{n,r}^{\flat}$ is the disjoint union of $\Lambda_{n,r}^{\flat}(1)$ and $\Lambda_{n,r}^{\flat}(2)$. We define
@@
$B_{1,r}:=\sum_{\sigma\in\Lambda_{n,r}^{\flat}(1)}\mathcal{E}_{r}(\sigma)^{t_{r}};\;B_{2,r}:=\sum_{\sigma\in\Lambda_{n,r}^{\flat}(2)}\mathcal{E}_{r}(\sigma)^{t_{r}}.$
In the following, we estimate $B_{1,r}$ and $B_{2,r}$ separately.
@@
$\sum_{\sigma\in T(\widetilde{\sigma})}\mathcal{E}_{r}(\sigma)^{t_{r}}=\sum_{\sigma\in T(\widetilde{\sigma})}(\mu(F_{\sigma})a_{\sigma_{L}^{r}})^{t_{r}}\geq\bigg(\sum_{\sigma\in T(\widetilde{\sigma})}\mu(F_{\sigma})a_{\sigma_{L}^{r}}\bigg)^{t_{r}}\geq\underline{a}^{rM_{r}t_{r}}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}.$
Let $H_{3,r}:=\underline{a}^{rM_{r}t_{r}}$. It follows that
(4.6) $B_{1,r}=\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}(1)}\sum_{\sigma\in T(\widetilde{\sigma})}\mathcal{E}_{r}(\sigma)^{t_{r}}\geq H_{3,r}\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}(1)}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}.$
For every $\sigma\in\Lambda_{n,r}^{\flat}(2)$, we have, $S(\sigma)\subseteq G_{1}(\sigma)$. This and Lemma 4.3 yield
(4.7) $B_{2,r}\geq\sum_{\sigma\in\Lambda_{n,r}^{\flat}(2)}H_{2,r}^{-1}\sum_{\widetilde{\sigma}\in S(\sigma)}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}=H_{2,r}^{-1}\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}(2)}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}.$